# EE-LLM: Large-Scale Training and Inference of Early-Exit Large Language Models with 3D Parallelism

### v2

@@ 4.2.2 Further performance optimizations
4.3 Advanced features
EE-LLM incorporates some advanced features that can potentially improve the training process, which are introduced below. We note that these are exploratory functionalities, and formal investigation of their practical benefits is left for future work.
4.3.1 Non-constant weights of early-exit losses
The weights of early-exit losses in the training objective of Eq. (1) can be changing rather than constant during the training process, just like the learning rate or other hyperparameters. Allowing the weights to be changing can offer a more fine-grained control of how gradients from multiple loss functions jointly impact the backbone and early/final output layers of the model. This functionality has been implemented in EE-LLM.
One concrete option that we offer is warm-up. With this option, early-exit loss weights start at small values, and gradually increase with training iterations until reaching the pre-specified maximum values. This approach has been adopted in prior works [31]. The idea is to encourage the deep neural network to primarily optimize for the full-model output quality from the beginning of the training process, while the skill of early exiting is gradually acquired with minor or no negative impact on the final outputs of the full model.
Another option is cool-down, which does the opposite and decreases early-exit loss weights during the training process. This option is inspired by prior works [39, 69] that leverage early-exit losses for the purpose of regularizing the training process of deep neural networks. Such “deep supervision” provided by early-exit losses can stabilize and accelerate convergence of training, and potentially improve the intermediate features learned by the neural network. As early-exit loss weights gradually decay, the regularization gets weaker, so that the network becomes more focused on its primary objective, namely the final-output quality.
4.3.2 Filling explicit bubbles with additional microbatches
To further leverage the explicit bubbles within the 1F1B pipeline schedule, i.e. the gray areas in Figure 5, we design and implement a novel approach of filling them with partial forward/backward computation of additional microbatches. This is primarily inspired by the idea from [51]: instead of designing a new, sophisticated pipeline schedule that has a lower bubble ratio, one may seek to fill the bubbles of an existing schedule with useful computation, which can lead to better resource utilization and faster training.
Our approach is visualized in Figure 5 and explained below. For notational simplicity, let us call the explicit bubbles between the warm-up and steady phases as Part 1, and the bubbles during the cool-down phase as Part 2. For each part, we fill them with some computation for $K$ additional microbatches. More concretely, for Part 1, the $i$-th inserted microbatch (where $i\in[K]$) goes through forward computation of the first $K+1-i$ pipeline stages, followed by backward computation of all visited early-exit losses; for Part 2, each inserted microbatch goes through forward computation of all stages, followed by backward computation of the final and early-exit losses (if any) only for the last few stages. In this way, each training iteration can process more data without any time overhead, as long as the number of inserted microbatches and the number of stages for partial forward/backward computation are chosen appropriately.
One thing to note is that this approach changes the optimization semantics. From the perspective of stochastic optimization, we can prove formally that with such additional computation and under certain conditions, one obtains an unbiased gradient estimate with reduced variance for the original training objective. We refer interested readers to Appendix B for more details about the methodology and analysis.
Figure 5: A visualization for our proposed method of filling pipeline bubbles with partial forward and backward computation of four additional microbatches, i.e. the ones annotated as P1, P2, P3 and P4.
5 Inference
This section first explains the major challenge faced by early-exit inference of LLMs in autoregressive generation tasks, namely the conflict between early exiting and KV caching, as well as a few recent attempts to resolve it. Then, we introduce our novel approach, which is based on a new type of pipeline parallelism. Throughout this section, we focus on the latency-oriented setting, with a batch size of 1 during sequence generation.
5.1 Main challenge: KV caching
Accelerating inference with early exiting is, in principle, orthogonal to and compatible with many other common techniques of acceleration [32], such as kernel fusion, FlashAttention [15], model quantization [85, 80], among others. For autoregressive generation tasks, one exception is KV caching, i.e. saving keys and values of all attention layers for previously generated tokens, which accelerates sequence generation by reducing redundant computation (though at the cost of higher memory usage). This is obviously in contradiction with vanilla early-exit inference: if the current token is generated via early exiting, then its KV caches in later layers are missing, which hinders the generation of future tokens that go beyond the exiting layer of the current token.
This challenge has been well recognized in the literature, and several approaches have been recently proposed to resolve it. One approach [19, 42, 61] is to copy the hidden states of the current token at the exiting layer to all later layers, which will be used to compute the keys and values at later attention layers. Despite its efficiency, this method obviously causes a deviation from the inference process that the model is trained to excel at, which can harm the output quality. Another solution [13] is to pre-specify the exiting layer for each token, while ensuring that KV missing in previous tokens will not hinder generation of later tokens; with this approach, the ability of token-wise adaptive selection of exits is inevitably lost. The third method [3, 70] is to store the hidden states of recent tokens that were generated with early exiting, and whenever KV missing happens, run a batch forward pass with the current and recent tokens to fulfill the KV caches. A variant of this approach, which we call KV recomputation, is implemented in EE-LLM; we defer its details to Section 6.3. Despite the batching effect of GPU computation, this approach still brings additional computational overhead, and might not even achieve any acceleration for inference on other hardware platforms, due to its high computational complexity.
5.2 A new solution: pipeline parallelism
We propose a novel solution that leverages a new type of pipeline parallelism during inference. The key idea is that, in the process of inference with multiple pipeline stages, the following two processes run in parallel whenever the model decides to do early exiting for the current token at a certain early exit:
  • The full-model forward pass of the current token is continued from the exiting layer, which fulfills the KV caches in all later layers.
See Figure 6 for a visualization of our proposed approach. Even though each token essentially goes through a forward pass of the full model, the computation after the exiting layer is parallelized with the computation of later tokens, which is how acceleration is achieved in this approach. It can be checked that the inference latency for generating one token at a certain exit matches exactly the time needed for the forward computation before returning an output at that exit, unless the selected exit is located in the middle of the first pipeline stage, in which case generation of the next token has to wait until the forward pass of the first stage for the current token is completed. Note that this is true not just in practice but also for the theoretical time complexity, without relying on the batching effect of GPU computation like KV recomputation does.
An implementation of this pipeline-based inference method is provided in EE-LLM. One potential limitation of the proposed method is that it requires multiple devices to facilitate pipeline parallelism, although parallelism within a single GPU or other device might be possible with more advanced implementation.
Figure 6: A comparison between the standard full-model inference (top), and our pipeline-based approach of early-exit inference (bottom). Numbers in the blocks represent the tokens within one generated sequence. Here, we consider the special case where each early exit is located at the end of a certain pipeline stage, just for simplicity of visualization.
6 Implementations
The implementation of EE-LLM is based on Megatron-LM [49], primarily extending Megatron-LM’s model architectures, pipeline scheduling, and inference service to support the training and inference of early-exit LLMs. We introduce each of these aspects in more details below.
6.1 Model architectures
6.2 Pipeline scheduling
We have adjusted the existing 1F1B schedule for early-exit LLMs, as shown in Figure 3. To fill implicit bubbles and reduce GPU memory overhead, lazy loss functions of early-exit layers are returned together with outputs of the backbone network during forward steps. These lazy functions are not actually called until their corresponding auxiliary losses (cf. Section 4.1.1) are calculated in the backward steps. For the method of filling explicit bubbles proposed in Section 4.3.2, we have inserted partial forward/backward computation of additional microbatches into warm-up and cool-down phases of the 1F1B schedule. The number of inserted microbatches and partial forward/backward stages can be automatically calculated through the user-specified (estimate of) ratio between backward and forward time.
6.3 Inference service
For inference with pipeline parallelism, i.e. the pipeline-based approach proposed in Section 5.2, we have re-implemented the forward process. With our implementation, the first pipeline stage will wait for an exit signal from the early/final exits of all subsequent stages after its forward computation is completed. Each subsequent stage will send an exit signal and the output token to the first stage, if there is an exit within the stage that satisfies the exit condition. Upon receiving the signal and generated token, the first stage will immediately start the forward pass for generating the next token. With this implementation, regardless of the early-exit layers’ positions in subsequent stages, the inference service can immediately generate a token whenever early exiting happens on some stage, without waiting for the completion of the entire stage (except for the first stage).
For inference without pipeline parallelism, we have implemented a mechanism of KV recomputation, which is a variant of synchronized parallel decoding proposed recently in [3]. In this approach, we maintain a list of the most recent tokens that have missing KV caches in deep layers due to early exiting. During each forward pass, we include these early-exit tokens in the current forward pass, which allows for direct recomputation of the KV caches for these tokens and thus avoids the issue of missing KV caches. Acceleration of sequence generation is still achieved, thanks to the batching effects of GPU computation. To avoid the endless accumulation of early-exit tokens, we enforce a full-model forward pass whenever the number of early-exit tokens reaches a pre-specified value.
7 Experiments
7.1 Training
In the following experiments, we empirically investigate the convergence of training early-exit models with EE-LLM, as well as the training efficiency of EE-LLM for early-exit LLMs up to an unprecedented scale of 30B. This scale is only limited by the hardware resources available to us, namely an 8-node cluster with 8 Nvidia A100-80GB GPUs in each node and hence 64 GPUs in total. We use a subset of the pre-training data [^5] https://github.com/alibaba/data-juicer/blob/main/configs/data_juicer_recipes/README.md provided by Data-Juicer [9], which consists of 800 billion tokens.
7.1.1 Convergence of training losses
First of all, we conduct experiments to verify the convergence of training losses when training early-exit LLMs with EE-LLM. We first consider a 1.3B GPT Transformer with 24 layers, add one minimalistic early-exit layer without layer normalization to the 1/4 depth and the other to the 1/2 depth, set their early-exit loss weights to 1/4 and 1/2 respectively (while the final-exit loss has a weight of 1), and tie all input and output embedding matrices. A standard LLM of the same GPT architecture is also trained using the same hyperparameters and pre-training data. We further train a 7B early-exit model with 32 layers using similar configurations, except that early-exit loss weights are set to 0.1 and 0.2, and all embedding matrices are untied. A 7B standard model is trained similarly. In all cases, we use the Adam optimizer [35] with hyperparameters $\beta_{1}=0.9,\beta_{2}=0.95,\epsilon=10^{-8}$, and a cosine schedule for the learning rate, with a maximum value of $3\times 10^{-4}$. The batch size and sequence length are both set to 2048.
Figure 7 shows the convergence of early-exit and final-exit training losses, i.e. negative log-likelihood of next-token prediction, for both standard and early-exit LLMs. All loss curves decay at a similar pace, and unsurprisingly, the early-exit losses are slightly higher than the final-exit loss for each model. Interestingly, the final-exit loss curve of each early-exit model is close to (or even slightly below) that of the standard model, suggesting that optimizing for early-exit losses might not hurt the final output of the full model in our setting. We also observe more spikes in the loss curves for the 7B model than for the 1.3B models, possibly due to the following reasons: (1) we choose to untie the embedding matrices and use smaller early-exit loss weights, both of which incur weaker regularization for the training process; (2) layer normalization, which is known to stabilize training of LLMs, is not included in the minimalistic early-exit layers of our 7B model.
(a) 1.3B, 24 layers
(b) 7B, 32 layers
Figure 7: Convergence of early-exit/final-exit training losses for our Transformer models. Each curve is annotated with the index of the Transformer layer that the corresponding exit is connected to.
7.1.2 Training efficiency
In this experiment, we investigate the training overhead of early exits. Starting with a standard GPT Transformer of size ranging from 1.3B to 30B, we increase the number of early exits from 0 to 3. Minimalistic early exits are added one by one to specific locations in the following order: (1) to the 1/4 depth; (2) to the 1/2 depth; (3) to the hidden states right before the first Transformer layer, which is always located in the first pipeline stage. Based on the performance optimizations proposed in Section 4.2, when an early exit is inserted into the middle of two layers located on two consecutive pipeline stages, we always add it to the beginning of the latter stage. We set the global batch size to 2048, microbatch size to 2 (for 1.3B and 7B models) or 1 (for 13B and 30B models), and sequence length to 2048. The degree of data parallelism is set to 4 and fixed throughout, while tensor/sequence and pipeline parallelism degrees can take various values.
Numerical results are illustrated in Figure 8, which matches the results from our analytical study in Section 4.2. In particular, training time per iteration (i.e. time for processing one data batch) increases with the number of added early exits, but at a slower rate when pipeline parallelism is enabled, thanks to the proposed utilization of implicit bubbles. Without pipeline parallelism, peak GPU memory increases with the number of early exits; on the other hand, with the pipeline parallelism degree set to 4, peak memory remains unchanged as early exits are added to middle pipeline stages, and only increases when the last early exit is added to the first stage, as predicted by our analysis in Section 4.2.
(a) 1.3B
(b) 7B
(c) 13B
(d) 30B
Figure 8: Training time per iteration and peak GPU memory vs. the number of added early exits, under various configurations of model sizes and parallelism degrees. Note that wall-clock time can be impacted by other workloads on the same GPU cluster when our experiments were conducted, which inevitably causes perturbations to our numerical results.
7.2 Inference
In the experiments below, we verify the effectiveness of our pipeline-based approach (with the number of pipeline stages set to 4) proposed in Section 5.2, and the method of KV recomputation explained in Section 6.3, for autoregressive generation with early exiting. Given that both methods generate the same output for the same prompt, we first investigate the downstream performance and early-exit speedup for the pipeline-based method alone, and then compare the inference speed of both methods. A server with 4 Nvidia A100-40GB GPUs is used for inference.
To examine the performance of inference with various sizes of models, we conduct experiments under different settings. One model that we use is an 1.3B early-exit model pre-trained from scratch with 300B tokens, and the other one is a 7B model pre-trained with 150B tokens. Each of them has two minimalistic early exits added to the 1/4 and 1/2 depth respectively; other configurations follow those in Section 7.1.1. Neither model has gone through any fine-tuning or alignment. Since designing the decoding method or early-exit condition is not the focus of our work, we simply adopt greedy decoding and a basic confidence-based decision rule for early exiting: at each early exit, if the maximum probability of next-token prediction is above a pre-defined threshold, then the model chooses the most likely token and moves on to generate the next one. The trade-off between output quality and speed of generation is then controlled by the threshold. Note that for a fair comparison, when the threshold is set to 1, early-exit layers are disabled, so that the computational complexity of a full-model forward pass matches that of a standard LLM without early exits; the inference latency with the threshold set to 1 is also used as the baseline for calculating (relative) speedup.
7.2.1 Output quality and speed of generation
We evaluate the downstream performance of our models with HELM [43] in four tasks. Among them, BoolQ [12] and TruthfulQA [44] are question-answering tasks, while XSUM [46] and CNN/DailyMail [64] are summarization tasks. We use EM (ratio of exact match with the correct answers) as the metric for the former, and ROUGE-L (a measure of similarity to the reference summaries) for the latter. Figure 9 demonstrates the relation between scores and inference speedup for pipeline-based inference with varying thresholds. Encouragingly, for many of these tasks, inference with early exiting can achieve 2×2$\times$2 × or higher speedup compared to full-model inference, with comparable or even better evaluation scores. Table 2 provides a concrete piece of text generated by our 7B early-exit model, showing that speedup can be achieved with minor or no impact on the resulted sequence. Table 3 goes one step further and demonstrates the confidence at each exit for each token, confirming the existence of easy tokens that can be predicted with high confidence at early exits, without wasting computational resources on later layers. It is worth noting that performance of early-exit inference might be improved in many ways, e.g. using early exits of different structures, other decoding mechanisms and exit conditions, or simply more extensive and sufficient pre-training. We also conjecture that early exiting can, in principle, bring higher speedup for larger models, as expressivity of early exits grows with the model capacity, which allows them to make confident and accurate predictions for a larger proportion of tokens.
(a) 1.3B
(b) 7B
Figure 9: Evaluation scores and relative speedup with our early-exit models. For each curve, the confidence threshold decreases from left to right.
Prompt
Artificial General Intelligence is
Full model
(Time: 1.42 s)
The term "artificial intelligence" was coined by John McCarthy in 1956.
$\text{Threshold}=0.8$
(Time: 1.17 s)
The term "artificial intelligence" was coined by John McCarthy in 1956.
$\text{Threshold}=0.4$
(Time: 0.89 s)
$\text{Threshold}=0.2$
(Time: 0.58 s)
Artificial General Intelligence is a branch of computer science that
[…]
Layer 8
Layer 16
Final layer
located
Be
Be
ij (0.992) ij (0.995) ij (0.996)
ing (0.999) ing (0.999) ing (0.999)
[…]
The
The
The
capital
capital
capital
of
of (0.832) of (0.860)
China
the
the
country
United
United
States (0.880) States (0.816) States (0.858)
is
is
is
Washington (0.877) Washington (0.992) Washington (0.989)
D
D
D
. (0.999) . (0.997) . (0.997)
C (0.969) C (0.967) C (0.934)
. (0.830) . (0.873) . (0.904)
7.2.2 A comparison between two inference methods
We compare the inference time of both pipeline-based method and KV recomputation. Since the pipeline-based method uses 4 GPUs for a pipeline parallelism (PP) degree of 4, we allow KV recomputation to use a tensor parallelism (TP) degree of 1 or 4 for a fair comparison. The empirical results on XSUM and CNN/DailyMail tasks are shown in Figure 10. We first notice that the pipeline-based approach outperforms KV recomputation with $\text{TP}=1$ for confidence thresholds smaller than 1 (i.e. when early exiting actually happens), despite the point-to-point communication overhead. Moreover, KV recomputation with $\text{TP}=4$ is faster than the pipeline-based approach; it is worth noting, however, that the small speedup from $\text{TP}=1$ to $\text{TP}=4$ for KV recomputation is only possible with high-end hardware like A100 GPUs connected by high-bandwidth communication. In many use cases of LLM inference, such hardware is not available, and pipeline parallelism is the only option for partitioning a model that is too large to fit into the memory of one single device. In sum, the proposed pipeline-based approach is a more practical option for LLM inference with model partitioning in a wide range of scenarios.
(a) XSUM
(b) CNN/DailyMail
Figure 10: A comparison of inference speed between the pipeline-based method (with 4 pipeline stages) and KV recomputation (with the degree of tensor parallelism set to 1 or 4) for our 7B early-exit model.
8 Related works
As introduced in Section 1, early exiting has been widely applied for accelerating inference of deep neural networks in the literature. In terms of early-exit Transformers in the NLP domain, the majority of prior works is focused on BERT [16] or other encoder-only models for classification tasks [45, 25, 94, 19, 63, 81, 62, 82, 42, 26]. Recent works have begun to study token-wise early exiting for accelerating inference of encoder-decoder or decoder-only LLMs in autoregressive sequence generation [62, 13, 3, 76]. While they are largely focused on designing inference mechanisms, the lack of support in prior works for training early-exit models with massive 3D parallelism inevitably posts an obstacle towards truly scaling up early-exit LLMs. Our work on EE-LLM is one important step towards eliminating this obstacle, and we also contribute a novel pipeline-based inference mechanism along the way.
Interestingly, early exiting has been used for many other purposes in the literature, besides accelerating inference. For example, early-exit losses can provide deep supervision and regularization for training deep networks with enhanced stability of convergence [69, 39], which motivates our implementation of the cool-down option for non-constant early-exit loss weights, as explained in Section 4.3.1. Other works use early exits for interpreting the intermediate features learned by each layer of a Transformer [37], for improving outputs of the final exit [22], or as the draft models in speculative decoding [33]. Early exiting has also found application in long-tailed classification [18], federated learning with heterogeneous clients [29, 93], token reduction [1], layer-wise training of deep networks [60, Section 5.3], among others. We believe that our progress in this work can be beneficial for these various purposes.
It is worth mentioning that there are other types of dynamic neural networks [24, 83] that facilitate conditional computation and elastic inference, such as layer skipping [6, 5, 79, 13, 17, 87, 78] and mixtures of experts [30, 66, 21]. Another line of work aims to accelerate LLM inference by designing new decoding methods rather than model architectures [40, 59, 91]. Each of these approaches has its own pros and cons, and some of them can be complemented by early exiting. A detailed comparison between these methods is beyond the scope of the current work.
9 Conclusions
We have introduced EE-LLM, a system for large-scale training and inference of early-exit LLMs with 3D parallelism. For training, we have presented how to execute backpropagation of the early-exit training objective across pipeline stages, various performance optimizations for minimizing the computational overhead compared to training standard LLMs, and some advanced features for more fine-grained control and optimization of the training process. For inference, we have introduced our design and implementation of two inference methods, one based on KV recomputation and the other based on a new type of pipeline parallelism, both of which are compatible with KV caching for autoregressive generation. Along the way, we have discovered some interesting chemical reaction between early exiting and pipeline parallelism, which is likely because it is both along the depth dimension that an early-exit model executes forward computation selectively and gets partitioned by pipeline parallelism. With EE-LLM, it now becomes possible to train and deploy early-exit LLMs that are as large as any standard LLM allowed by Megatron-LM, given the same amount of computational resources. We hope that EE-LLM will be a helpful tool for future study and applications of early-exit LLMs at larger scales.
At the time of writing this article, we are continuously updating our codebase with more thorough experiments and more advanced functionalities. It is worth noting that many ideas presented in this work can be generalized to broader settings, e.g. deep neural networks other than decoder-only GPT Transformers, frameworks other than Megatron-LM, and hardware platforms other than GPU clusters; it would be exciting to see further developments in these aspects.
Appendix A Analysis of training efficiency
In this section, we derive formulas for the training time per iteration and peak GPU memory usage across pipeline stages under general configurations of early exits, in terms of the quantities listed in Table 4. Some of these quantities, especially $\{f_{\mathsf{o}},b_{\mathsf{o}},m_{\mathsf{o}},m^{\dagger}_{\mathsf{o}}\}$, can be further calculated analytically with lower-level quantities such as sequence length, microbatch size, hidden size, vocabulary size and others [57, 49, 36], or estimated empirically by profiling in practice.
@@ Appendix A Analysis of training efficiency
A.1 Training time per iteration
@@
We start by recalling the simplified analysis in prior works for the standard 1F1B pipeline schedule [49], which assumes that all stages have the same forward time $f$ and backward time $b$ for one microbatch. Recall from Figure 3 that the critical path of one training iteration consists of three parts:
@@
A.2 Peak GPU memory
Consider the GPU memory for one specific pipeline stage, say Stage $i$. Recall that GPU memory usage during training is mainly composed of memory for model parameters, gradients, optimizer states, and intermediate activations [57, 36]. Among them, the memory for gradients and optimizer states (for many common optimizers, such as SGD with momentum or Adam) is proportional to the number of model parameters. Therefore, we assume that
@@ A.2 Peak GPU memory
The memory for intermediate activations can be similarly divided into components corresponding to $\mathsf{BB}$, $\mathsf{IN}$, $\mathsf{FE}$ and $\mathsf{EE}$. Note that according to Section 4.2.2, the peak memory usage by activations within one early-exit layer is only $m^{\dagger}_{\mathsf{EE}}$ regardless of the number of in-flight microbatches $P+1-i$. Thus one has
@@
Appendix B Supplementary materials for Section 4.3.2
This section provides more details about the method, proposed in Section 4.3.2, of filling pipeline bubbles with partial forward/backward computation of additional microbatches.
B.1 Methodology

### v3

@@ A.2 Further performance optimizations
Numerical examples.
A.3 Theoretical analysis for general cases
In the following, we derive formulas for the training time per iteration and peak GPU memory usage across pipeline stages under general configurations of early exits, in terms of the quantities listed in Table 2. Some of these quantities, especially $\{f_{\mathsf{o}},b_{\mathsf{o}},m_{\mathsf{o}},m^{\dagger}_{\mathsf{o}}\}$, can be further calculated analytically with lower-level quantities such as sequence length, microbatch size, hidden size, vocabulary size and others (Rajbhandari et al., 2019; Narayanan et al., 2021b; Korthikanti et al., 2022), or estimated empirically by profiling in practice.
@@ A.3 Theoretical analysis for general cases
A.3.1 Training time per iteration
@@
We start by recalling the simplified analysis in prior works for the standard 1F1B pipeline schedule (Narayanan et al., 2021b), which assumes that all stages have the same forward time $f$ and backward time $b$ for one microbatch. Recall from Figure 3 that the critical path of one training iteration consists of three parts:
@@
A.3.2 Peak GPU memory
Consider the GPU memory for one specific pipeline stage, say Stage $i$. Recall that GPU memory usage during training is mainly composed of memory for model parameters, gradients, optimizer states, and intermediate activations (Rajbhandari et al., 2019; Korthikanti et al., 2022). Among them, the memory for gradients and optimizer states (for many common optimizers, such as SGD with momentum or Adam) is proportional to the number of model parameters. Therefore, we assume that
@@ A.3.2 Peak GPU memory
The memory for intermediate activations can be similarly divided into components corresponding to $\mathsf{BB}$, $\mathsf{IN}$, $\mathsf{FE}$ and $\mathsf{EE}$. Note that according to Section A.2, the peak memory usage by activations within one early-exit layer is only $m^{\dagger}_{\mathsf{EE}}$ regardless of the number of in-flight microbatches $P+1-i$. Thus one has
@@
Appendix B Additional experiments
B.1 A comparison between two inference methods
Since both pipeline-based method and KV recomputation generate the same output for the same prompt, here we only compare the inference latency of both methods. Given that the pipeline-based method uses 4 GPUs for a pipeline parallelism (PP) degree of 4, we allow KV recomputation to use a tensor parallelism (TP) degree of 1 or 4 for a fair comparison. The empirical results on XSUM and CNN/DailyMail tasks are shown in Figure 10. We first notice that the pipeline-based approach outperforms KV recomputation with $\text{TP}=1$ for confidence thresholds smaller than 1 (i.e. when early exiting actually happens), despite the point-to-point communication overhead. Moreover, KV recomputation with $\text{TP}=4$ is faster than the pipeline-based approach; it is worth noting, however, that the small speedup from $\text{TP}=1$ to $\text{TP}=4$ for KV recomputation is only possible with high-end hardware like A100 GPUs connected by high-bandwidth communication. In many use cases of LLM inference, such hardware is not available, and pipeline parallelism is the only option for partitioning a model that is too large to fit into the memory of one single device. In sum, the proposed pipeline-based approach is a more practical option for LLM inference with model partitioning in a wide range of scenarios.
(a) XSUM
(b) CNN/DailyMail
Figure 10: Inference latencies of the pipeline-based method and KV recomputation for our 7B early-exit model.
B.2 Examples of generated texts
Example texts generated by our early-exit models are presented in Tables 3 and 4.
Prompt Artificial General Intelligence is
Full model
(Time: 1.42 s)
The term ”artificial intelligence” was coined by John McCarthy in 1956.
$\text{Threshold}=0.8$
(Time: 1.17 s)
The term ”artificial intelligence” was coined by John McCarthy in 1956.
$\text{Threshold}=0.4$
(Time: 0.89 s)
$\text{Threshold}=0.2$
(Time: 0.58 s)
Artificial General Intelligence is a branch of computer science that
[…]
Layer 8 Layer 16 Final layer
located (0.116) Be (0.202) Be (0.609)
ij (0.992) ij (0.995) ij (0.996)
ing (0.999) ing (0.999) ing (0.999)
, (0.370) . (0.535) . (0.503)
[…]
The (0.126) The (0.150) The (0.158)
capital (0.074) capital (0.125) capital (0.102)
of (0.757) of (0.832) of (0.860)
China (0.523) the (0.094) the (0.115)
country (0.146) United (0.438) United (0.436)
States (0.880) States (0.816) States (0.858)
is (0.733) is (0.617) is (0.687)
Washington (0.877) Washington (0.992) Washington (0.989)
, (0.443) , (0.394) , (0.462)
D (0.571) D (0.728) D (0.686)
. (0.999) . (0.997) . (0.997)
C (0.969) C (0.967) C (0.934)
. (0.830) . (0.873) . (0.904)
B.3 Training with different early-exit configurations
As mentioned in Section 2, EE-LLM supports a wide variety of early-exit configurations, such as numbers, locations, structures and training loss weights of early exits. Here, we demonstrate the training of two additional 7B early-exit models. The first model has the same configurations as those of the 7B early-exit model considered in Section 5.1, except that each early-exit layer contains layer normalization and an MLP, on top of the output embedding matrix. The second model has three early exits located at Layer 4, 12 and 20, respectively, all containing layer normalization, with early-exit loss weights set to 0.1, 0.2 and 0.3. In addition, all embedding matrices are tied. Our empirical results, as shown in Figure 11, confirm the convergence of training losses for both models. It would be interesting future work to compare more thoroughly different early-exit configurations for both training and inference.
Figure 11: Convergence of early-exit/final-exit training losses for two additional models with different configurations. Each curve is annotated with the index of the Transformer layer that the corresponding exit is connected to. The missing parts in the curves of the first figure were caused by technical issues with wandb (https://wandb.ai) when the experiments were running.
Appendix C Advanced features
This section presents more details about the advanced features introduced in Section 3.3 for training early-exit LLMs.
C.1 Non-constant early-exit loss weights
The weights of early-exit losses in the training objective of Eq. (1) can be changing rather than constant during the training process, just like the learning rate or other hyperparameters. Allowing the weights to change can offer more fine-grained control of how gradients from multiple losses jointly impact the backbone and early/final output layers of the model.
One concrete option that we offer in EE-LLM is warm-up. With this option, early-exit loss weights start at small values, and gradually increase with training iterations until reaching the pre-specified maximum values. This approach has been adopted in prior works (Kaya et al., 2019). The idea is to encourage the deep neural network to primarily optimize for the full-model output quality from the beginning of the training process, while the skill of early exiting is gradually acquired with minor or no negative impact on the final outputs of the full model.
Another option is cool-down, which does the opposite and decreases early-exit loss weights during the training process. This option is inspired by prior works (Lee et al., 2014; Szegedy et al., 2015) that leverage early-exit losses for the purpose of regularizing the training process of deep neural networks. Such “deep supervision” provided by early-exit losses can stabilize and accelerate the convergence of training, and potentially improve the intermediate features learned by the neural network. As early-exit loss weights gradually decay, the regularization gets weaker, so that the network becomes more focused on its primary objective, namely the final-output quality.
C.2 Filling explicit bubbles
To further leverage the explicit bubbles within the 1F1B pipeline schedule, i.e. the light gray areas in Figure 4, we design and implement a novel approach of filling them with partial forward/backward computation of additional microbatches. See Figure 4 for a visualization. This is primarily inspired by the idea from (Osawa et al., 2023): instead of designing a new, sophisticated schedule that has a lower bubble ratio, one may seek to fill the bubbles of an existing schedule with useful computation, which leads to better resource utilization and faster training. Note that our method changes the optimization semantics. From the perspective of stochastic optimization, we can prove formally that with such additional computation and under certain conditions, one obtains an unbiased gradient estimate with reduced variance for the original training objective. The remaining of this section is dedicated to details of the methodology and analysis.
C.2.1 Methodology
Our approach is explained below. For notational simplicity, let us call the explicit bubbles between the warm-up and steady phases as Part 1, and the bubbles during the cool-down phase as Part 2. For each part, we fill them with some computation for $K$ additional microbatches. More concretely, for Part 1, the $i$-th inserted microbatch (where $i\in[K]$) goes through forward computation of the first $K+1-i$ pipeline stages, followed by backward computation of all visited early-exit losses; for Part 2, each inserted microbatch goes through forward computation of all stages, followed by backward computation of the final and early-exit losses (if any) only for the last few stages. In this way, each training iteration can process more data without any time overhead, as long as the number of inserted microbatches and the number of stages for partial forward/backward computation are chosen appropriately.